Click Here to receive Service Notes automatically via email.
Date | Description |
---|---|
03/25/2004 | Campus Pipeline was unavailable from approximately 9:20 - 9:30 am on 3/25/2004. IT Services is continuing to work with Campus Pipeline support. |
03/25/2004 | Campus Pipeline was unavailable from approximately 6:20 - 6:40 pm on 3/24/2004. IT Services is working with Campus Pipeline support on this issue. |
03/19/2004 | The Truman Home Page was mistakenly set back to an old version from about 4:20 - 8:05 am on 3/19/04. This affected any pages on the site www.truman.edu. The proper home page has now been restored. |
03/16/2004 | The slow Internet problem that began at ~ 11:15 this morning has been resolved. Several persons computer systems appear to have been hacked and were being used to lanch a denial of service attack against an outside (non-Truman) computer system. The problem computer systems have been disconnected from the campus network. |
03/10/2004 | Internet connectivity was lost briefly between 3:20 and 3:35. Service has been restored. We are investigating the cause. |
03/10/2004 | The wireless network in McClain was down briefly due to additional network drops being added in the help desk area. The network has been restored. |
03/10/2004 | Campus Pipeline was unavailable from approximately 11:30-11:40 a.m. for an upgrade of the Antivirus Software on the server. |
03/10/2004 | Campus Pipeline was briefly unavailable around 7:30 a.m. on 3/10/04 to test monitoring software. |
03/10/2004 | The administrative web server argon will be unavailable on and off from 7:20 a.m. until approximately 7:50 a.m. on 3/10/04 IT Services is performing an upgrade to the antivirus software, which requires several restarts of the server to complete. |
03/03/2004 | Faculty and staff W: and Y: network drives on the TS1 server were unavailable for several minutes during the lunch hour due to a problem with the server. |
03/01/2004 | Faculty and Staff network drives were briefly unavailable at ~11:10 AM due to a server problem. |
03/01/2004 | Exchange email services were briefly unavailable this morning at ~ 9:30 AM due to a disk space problem on the server. |
03/01/2004 | Email sending was down on Sunday evening for ~ 2 hrs. Service was restored at 11:00 pm We are investigating the cause of the problem. |
02/26/2004 | Email sending outage from 11:45-11:55 today. Service has been restored and the problem is under investigation. |
02/26/2004 | Internet access has been reported as slow. We are currently investigating. UPDATE 10:15AM Service has been restored. MyDoom virus floods ran the PacketShaper out of memory and forced a reboot. |
02/25/2004 | Internet access was disrupted this afternoon by several computers generating millions of myDoom virus packets. The infected computers are being disconnected from the network. |
02/23/2004 | Email sending was down on Sunday afternoon from ~ noon until 4:30 PM. We are investigating the cause of the problem. |
02/23/2004 | Network Y: drives were unavailable for several hours on Sunday due to a server problem. We are investigating the cause of the problem. |
02/19/2004 | The faculty and staff w: and Y: drives were briefly unavailable at 2 PM today due to a problem with the system virus protection. |
02/19/2004 | During tonights monthly IT maintenance window the Administrative Office V: drive will first be backed up and then moved to a new server. This change should be mostly invisible to network users since V: is automatically mapped by the network logon script. If you have access to V: and you leave yourself logged on to the network Thursday night you will need to log off Friday morning and then log back on to get the new V: drive. |
02/19/2004 | A defective network stack interface card caused several network related problems in the VH2000 computer lab and all other locations in the center core of Violette Hall this morning. The failed component has been replaced. |
02/19/2004 | Faculty and staff W: and Y: drives were briefly unavailable early this morning due to a problem with network connectivity after a system restart. The problem has been corrected. |
02/20/2004 | During the regularly scheduled maintenance window Thursday night, the Web Services team will be moving MS SQL services to a new server. The migration should be completed in approximately three hours. This will affect several database-driven web applications, such as the Admissions application, the ITS website and others. During the maintenance window these websites/web applications will be intermittently unavailable. |
02/18/2004 | Campus Pipeline was unavailable from approximately 1:00 pm - 2:00 pm on Tuesday, February 17. The cause of the outage is being investigated by Campus Pipeline support. |
02/16/2004 | 1 PM Update - MOREnet has corrected the problem with slow Internet and Mobius access. However, they said what they did was a "patch" and we may loose services again this afternoon. They are working on a plan to actually fix the problem later tonight. |